OpenStudy (credmond):

Can someone show me step by step how to do this problem? I'd really appreciate it! :) Let P(-3,1) and Q(5,6) be two points in the coordinate plane. Find an equation for the circle for which the segment PQ is a diameter.

OpenStudy (anonymous):

What is the distance between these points? That will give us the diameter of the circle, which we need.

OpenStudy (anonymous):

We also need to know the midpoint, because that is the center.

OpenStudy (anonymous):

Do you know how to do these things?

OpenStudy (credmond):

I think so! For the distance, I got sqrt89 and for the midpoint I got (1, 7/2)

OpenStudy (credmond):

is that right so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay so do you know the equation of a circle then?

OpenStudy (credmond):

I understand the first part of the equation, but for the r^2 portion, do I just take the square root of 89?

OpenStudy (credmond):

"take" meaning use the square root of 89

OpenStudy (credmond):

so... (x-1)^2 + (y-7/2)^2= sqrt89

OpenStudy (anonymous):

well, we need the radius, not diameter.

OpenStudy (credmond):

ohhh. got you.

OpenStudy (anonymous):

and it needs to be squared

OpenStudy (credmond):

so the radius is half of 89? or?

OpenStudy (credmond):

ohhhh wait never mind

OpenStudy (credmond):

so its sqrt89/2 squared?

OpenStudy (anonymous):

yeah, but when you square a sqrt it isn't a sqrt anymore

OpenStudy (credmond):

right!

OpenStudy (credmond):

so the whole equation is (x-1)^2 + (y+7/2)^2 = 89/4 ?

OpenStudy (anonymous):

that looks right

OpenStudy (credmond):

Woohoo! Thank you so much! I really appreciate it! :)